OpenStudy (anonymous):

How many ways can five members of the 100-member United States Senate be chosen to be put on a committee?

OpenStudy (anonymous):

100*99*98*97*96

OpenStudy (anonymous):

100!/(100-5)!

OpenStudy (anonymous):

or else use the method i told u on the movie question

OpenStudy (anonymous):

\[100P_5\] \[= \frac{100!}{(100-5)!}\]

OpenStudy (kropot72):

The order of the 5 selected is not important. 100C5 = \[\left(\begin{matrix}100 \\5\end{matrix}\right)=\frac{100!}{5!95!}\]